OpenStudy (anonymous):

Find a cubic function with the given zeros. Sqrt(2), -Sqrt(2), -2

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so if x = -2 is a zero then x + 2 is a factor so P(x) = (x + 2) .... now for the other factors (x - sqrt(2)) and (x + sqrt(2)) the product of these binomials is the difference of 2 squares so its x^2 - 2 if you think about it x^2 - 2 = 0 then x^2 = 2 and hence the zeros by taking the square root.. so now you know P(x) = (x +2)(x^2 - 2) just distribute for your answer. hope it helped